A car is traveling around a horizontal circular track with radius r- 280 m angle θ~ 26. above the x axis, and the angle θΒ-51, below the x axis. at a constant speed v 19 m/s as shown. The 1) What is the magnitude of the cars acceleration? 1.28 m/s2 Submit Your submissions:1.28 Computed value: 1.28 Smitted: Tuesday, January 29 at 9:21 PM Feedback: Correct 2) What is the x component of the cars acceleration when it is at point A m/s2 Submit 3) What is the y component of the cars acceleration when it is at point A m/s Submit
4) What is the x component of the cars acceleration when it is at point B s Submit 5) What is the y component of the cars acceleration when it is at point B s Submit

Please help with whatever you can! Thanks
Answer #1

1)

acceleration = v2/r = (19)2/(280) = 1.28 m/s2

2)

x component of acceleration is

ax = (1.28)*(cos 26) = 1.15 m/s2 (In the negative x direction)

3)

y component of acceleration is

ay = 1.28*(sin 26) = 0.56 m/s2

4)

x component = (1.28)(cos 51) = 0.805 m/s2 (In the negative x direction )

5)

y component = (1.28)(sin 51) = 0.994 m/s2